### v2

# Ensemble Debiasing Across Class and Sample Levels for Fairer Prompting Accuracy

Abstract
Language models are strong few-shot learners and achieve good overall accuracy in text classification tasks, masking the fact that their results suffer from great class accuracy imbalance. We believe that the pursuit of overall accuracy should not come from enriching the strong classes, but from raising up the weak ones. To address the imbalance, we propose a Heaviside step function based ensemble debiasing method, which enables flexible rectifications of in-context learned class probabilities at both class and sample levels. Evaluations with Llama-2-13B on seven text classification benchmarks show that our approach achieves state-of-the-art overall accuracy gains with balanced class accuracies. More importantly, we perform analyses on the resulted probability correction scheme, showing that sample-level corrections are necessary to elevate weak classes. Due to effectively correcting weak classes, our method also brings significant performance gains to a larger model variant, Llama-2-70B, especially on a biomedical domain task, further demonstrating the necessity of ensemble debiasing at both levels.
1 Introduction
Language models are good few-shot learners, but prompting can result in inevitable class accuracy imbalance. A language model can perform significantly better on certain classes compared to others, leading to biased performances across different classes (illustrated in Figure 1). This issue is particularly relevant in tasks like classification, or any scenario where the model must distinguish between multiple categories. For example, on DBpedia ontology classification [1], the overall accuracy for 14 classes can reach 88% when prompting Llama-2-13B [2], but class Nature’s accuracy is 28%. Broadly, for open-ended generation tasks, when treating each token in the vocabulary as a distinct class, generating the most likely next token can be alternatively viewed as classifying over the entire vocabulary or soft labels [3, 4], also susceptible to class accuracy imbalance. Therefore, accuracy imbalance is prevalent and affecting users of language model applications, where classification is the most basic case manifesting such imbalances.
Figure 1: The accuracy imbalance and what we can achieve with debiasing
In text classifications, the prevalence of class accuracy imbalance is sometimes hidden by a single or a dominant evaluation metric, such as overall accuracy. Especially for relatively large models, the fact that they excel on benchmarks which emphasize overall performances may hide a poor performance in some classes. To enable a focused evaluation of whether language model performances are balanced across classes, the COBias metric was proposed [5] to dedicate evaluating pairwise class accuracy differences.
The class accuracy imbalance issue is not particularly easy to solve from the root. Its causes come from within the model architecture. In details, language models rely on priors learned during training to make predictions. However, built on self-attention, the models are prone to the copying mechanism of the induction heads [6, 7]. These heads predict repeated sequences, which can over-capture surface-level patterns in the data, leading to biased priors, and then biased predictions. Therefore, if certain classes or concepts were more frequently emphasized in the training data, the model may disproportionately favor them [8]. Although a systemic change at model level or selecting high-quality data may help, it can be inefficient.
Besides what is rooted in the model and data, the way prompts are phrased can further bias the model towards certain classes [9, 10, 11]. In few-shot prompting, imbalanced class accuracy can be exacerbated because of underrepresented classes in the demonstrations of a prompt. Nevertheless, it only helps to a limited extent by making all classes well-represented in the demonstrations [5]. In addition, Chain-of-Thought style outputs may facilitate self-correcting responses without relying on prompt engineering [12], but the outcome is guided and restricted by reward signals. Lastly, although certain types of biases, such as gender bias or truthfulness, can be traced back to Transformer layers and intervened at inference time [13, 14, 15], interventions can be inefficient to mitigate the accuracy bias, especially when it could be better treated without altering the model’s internal structure.
Therefore, to effectively and efficiently reduce class accuracy imbalances, post-hoc optimization methods are introduced. As such, post-hoc probability correction complements training and prompting strategies to prevent the model from overly relying on learned patterns, and to incorporate necessary information in downstream tasks. This type of correction involves correcting In-context learned (ICL) output probabilities in inference time by combinatorially optimizing a relatively small mathematical model. These direct rectifications are useful, because they can directly correct outputs, no matter what biases are encoded in the model prior, or what prompts are used.
Related Work. Prior post-hoc correction work has investigated re-weighting ICL output probabilities through combinatorial optimization addressing COBias[5, 16] or calibration addressing biased predictions [17, 8]. The former uses labeled instances for optimization while the latter does not. Both lines of work advance overall classification accuracy with low costs and the flexibility of no language model parameter updates. The performance difference is that the combinatorial optimization works achieve better balanced class performances while improving overall accuracy. We follow the combinatorial optimization perspective for our ensemble debiasing.
For a background on combinatorial optimization, it provides a mathematical framework for maximizing or minimizing an objective function of decision variables, subject to constraints on some or all of the variables, involving inequalities, equalities, and integrality restrictions [18]. This framework essentially transforms a rich variety of tasks into discrete optimization models, which is robust and remarkably versatile. It solves real-world problems such as planning, scheduling, etc. in operations research, and has diverse applications in NLP, from structured prediction such as named entity recognition and dependency parsing, to reading comprehension, then to grammatical error correction, and to align LLM outputs with human preferences [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. The core use of the framework is to translate a problem into a mathematical statement of one of the forms of linear mixed-integer programming problem (MIP), integer programming problem (IP), nonlinear integer programming problem (NIP), or other types [32, 33, 34, 30]. Solving combinatorial optimization models is not an easy task. Among which, NIP is highly challenging, because the NIP model can contain discontinuous functions and is inherently non-differentiable. Lin and You [5, 16] proposed to address the search difficulty with metaheuristics, which we follow (more discussions in Appendix A). Back to the combinatorial optimization based debiasing methods, their drawbacks lie in the single consideration at either class level [5] or sample level [16]. In [5], any sample gets re-weighted by the same set of per-class correction weights, making it hard to capture sample-level variances, which is particularly needed for weak classes. In [16], corrections are made at sample level by assigning per-sample per-class correction weights that are computed by membership functions, however, capturing too much nuance may instead over-corrects strong classes.
To overcome the insufficiency in previous works, we propose a meticulous post-hoc correction framework DCS (Debiasing at Class and Sample levels). DCS is an organic integration of broad and fine-grained corrections across class and sample levels, pushing forward fairer prompting accuracy for LLMs. DCS automatically detects which classes need corrections that account for sample-level differences and which do not. In the following, we first introduce the Heaviside step function based ensembling strategy for DCS (Section 2) and the overall framework (Section 3), followed by quantitative evaluations and in-depth analyses (Section 4). In Section 5, we conclude with prospective applications and open problems.
Our contributions are summarized as below:
  • Post-hoc debiasing unifies class and sample level corrections. We ensemble weight correction and membership correction to enable rectifications of class probabilities at both class and sample levels, to enhance the performance of LLMs directly from their outputs.
  • SoTA scores on overall accuracy while maintaining low COBias DCS can achieve leading overall accuracy improvements while good COBias reduction across seven benchmark text classification tasks compared to the previous SoTA.
  • Optimal solutions need weight correction to elevate weak classes. Quantitative analyses show that the average ratio of classes using membership correction to weight correction is between 0 and 1, where most classes apply weight corrections but lowest-accuracy (weakest) classes more often apply membership corrections, suggesting the necessity of keeping membership correction as an option and letting the optimization process determine if it is needed for any particular class.
2 Using Heaviside Step Function to Ensemble Class and Sample Level Debiasing
We introduce a robust ensemble debiasing framework across class and sample levels, which is targeted specifically for balancing class accuracy while improving overall accuracy over individual methods. To achieve ensembling, a Heaviside step function seamlessly unifies two unique ICL output correction methods, bringing together the advantages of weight coefficient and fuzzy rule based corrections. This unified framework captures the intricacies of ICL output class probabilities with greater precision, reaching solutions that elevate weak classes while maintaining the strong ones. List of notations are presented in Appendix B.
2.1 The Mapping Function for ICL Probability Correction
As mentioned, a common method for correcting probabilities involves multiplying each class probability by a coefficient [5]. This is a class-level correction. For an $N$-class task, we use a set of weights to adjust each dimension of the ICL output $N$-dimensional probabilities without considering the variations in the initial probabilities across different instances. In contrast, a more refined correction method employs membership functions from fuzzy logic [16], where a mapping function is applied to adjust each dimension. Different classes use different mapping functions. As a result, class probabilities of different instances undergo varying degrees of transformation - some are adjusted significantly, others minimally, and some remain unchanged, reflecting a finer, sample-level correction approach.
[…]
$H(x)=\begin{cases}1,&x\geq 0\\ 0,&\text{otherwise}\end{cases}$ (1)
Let $\xi_{i}$ denote the selection variable for class $i$, each $\xi_{i}$ will be optimized to choose a most suitable correction from $\Psi$. We conventionally note when $\xi_{i}$ is optimized to be one of the indices from $\{1,...,D_{F}\}$, the i-th class probability of each instance gets a personalized adjustment by a membership function determined by $\xi_{i}$; when $\xi_{i}$ is one of the indices from $\{D_{F}+1,...,D_{F}+D_{W}\}$, the i-th class probability of any input instance is corrected by a weight coefficient determined by $\xi_{i}$. Consequently, when $\xi_{i}=k$ and $k\leq D_{F}$, the triangular membership function (following [16]) based correction function $\mu_{k}$ is selected, the updated probability is computed by:
$\mu_{k}(p_{mi})=\left\{\begin{aligned} &0,&&p_{mi}\leq a_{k}\\ &\frac{p_{mi}-a_{k}}{b_{k}-a_{k}},&&a_{k}<p_{mi}\leq b_{k}\\ &\frac{c_{k}-p_{mi}}{c_{k}-b_{k}},&&b_{k}<p_{mi}\leq c_{k}\\ &0,&&\text{otherwise}\end{aligned}\right.$ (2)
Special cases: when $a_{k}=b_{k}=0,\mu_{k}(p_{mi})=\frac{c_{k}-p_{mi}}{c_{k}}$ for $0\leq p_{mi}\leq c_{k}$ and 0 for $p_{mi}\geq c_{k}$; when $b_{k}=c_{k}=1,\mu_{k}(p_{mi})=\frac{p_{mi}-a_{k}}{1-a_{k}}$ for $a_{k}\leq p_{mi}\leq 1$ and 0 for $p_{mi}\leq a_{k}$.
Otherwise, the weight correction function $\omega_{k}$ is selected, and the updated probability is obtained by:
$\omega_{k}(p_{mi})=\frac{k-D_{F}}{D_{W}}p_{mi}$ (3)
Therefore, the per-sample per-class probability $p_{mi}$ is precisely updated as follows.
$p^{\prime}_{mi}=f_{i}(p_{mi})=\mu_{\xi_{i}}(p_{mi})\cdot H(D_{F}-\xi_{i})+\omega_{\xi_{i}}(p_{mi})\cdot H(\xi_{i}-D_{F}-1)$ (4)
This results in a more accurate correction for the ICL output probability, due to applying flexible fixes at either the broad level or the fine-grained level.
2.2 Mathematical Model
The mathematical model used to drive the selection is based on the nonlinear integer programming model in [5] with accuracy-maximizing and COBias-minimizing objectives. The combination of weight and membership corrections that optimizes the model’s objective function will be selected. In short, the model is:
$\min Z$ $=1-Z^{\text{Acc}}(\xi)+\beta Z^{\text{COBias}}(\xi)+\tau Z^{\text{PMI}}(\xi)$
s.t. $\boldsymbol{\xi}=(\xi_{1},\dots,\xi_{N}),\xi_{i}\in\{1,\dots,D_{F}+D_{W}\}$ (5)
where $1-Z^{\text{Acc}},Z^{\text{COBias}},Z^{\text{PMI}}$ respectively corresponds to the three terms of Eq. 5 in [5]. During optimization, at the beginning of each simulated annealing iteration, candidate selection $\xi_{i}$’s are generated and the correction functions $f_{i}$’s correct each sample in the optimization set, yielding corrected predictions $\{\hat{y}_{m}\}_{m=1}^{M}$. Each of the individual term in $Z$ is essentially an evaluation score computed using the corrected predictions and ground-truth labels $\{y_{m}\}_{m=1}^{M}$. If the new $Z$ is smaller than the current $Z$, the candidate selection is accepted and the best selection is updated; otherwise, the algorithm decides whether to accept the new selection based on the acceptance probability. Iterations repeat until a stopping criterion is met. The optimization process will be elaborated in Section 3.
Model parameters. This is a budget-friendly model with a total number of parameters (here we refer to search space size) of $N(D_{F}+D_{W})$.
3 Simulated Annealing for Solving DCS
Due to the non-differentiable nature of the math model, we turn to heuristics and employ the simulated annealing (SA) algorithm to solve the DCS model. The SA algorithm was proposed by [35, 36]. The forerunners successfully introduced the concept of annealing into combinatorial optimization, inspired by the annealing process of solid materials. It starts at an initial temperature and cools down at a specified rate. As the temperature gradually decreases, the algorithm leverages probabilistic jumps to explore the solution space randomly, searching for the global optimum of the objective function, so it can probabilistically escape local optima and eventually converge to the global optimum [37]. The optimization of DCS model is a search problem in multidimensional space, where each solution makes a potential probability correction scheme. Next, we uncover the overall framework that apply SA to solve our model.
3.1 Overall Framework
The overall solution framework for DCS is depicted in Fig. 2. The combinatorial optimization relies on a labeled set of these initial class probabilities and their ground-truth labels. The labeled set can be simply taken from the full or a subset of the classification task’s training set. After ICL output class probabilities are obtained, the optimization process starts with an initial solution. The iteration repeats until either a minimum temperature is reached or a maximum number of iterations is reached, and we obtain the optimal correction functions. In actual inference, the selected functions directly rectify a test instance’s ICL output class probabilities and result in corrected predictions.
Figure 2: Overall framework
3.2 Initial Solution
[…]
3.3 Neighborhood Solution
It is necessary to generate a new solution in the neighborhood of the current solution. If the current solution is $\xi_{j}=k_{j},\text{ for class }j=1,\dots,N$, we can randomly select a class, for example, $j^{\prime}$ whose current solution is $\xi_{j^{\prime}}=k_{j^{\prime}}$, and perturb it by sampling $k_{j^{\prime}}^{\text{New}}$ from the set $\mathbb{D}\setminus k_{j^{\prime}}$ as the new value for the variable $\xi_{j^{\prime}}$. That is, $\xi_{j^{\prime}}=k_{j^{\prime}}^{\text{New}}$. In this way, we obtain a neighboring solution: $\boldsymbol{\xi}^{\text{New}}=(\xi_{1}=k_{1},\dots,\xi_{j^{\prime}}=k_{j^{\prime}}^{\text{New}},\dots,\xi_{N}=k_{N})$. The transition from the current solution at a given temperature to the new solution is a Markov chain [35, 36], and the transition probability is determined by the Metropolis criterion $P_{\text{accept}}=e^{-\Delta Z/T}$. If the new solution is better ($\Delta Z<0$), accept the new solution; if the new solution is worse ($\Delta Z\geq 0$) and a uniform random number $r\sim U(0,1)$ is smaller than $P_{\text{accept}}$, namely $r<P_{\text{accept}}$, the new solution is also accepted.
3.4 Annealing Schedule
An annealing schedule is characterized by the initial value of the temperature, the cooling rate, the inner-loop criterion, and the outer-loop criterion.
Initial Temperature:The initial temperature is chosen so that the acceptance probability is relatively close to one. The higher the initial temperature, the more possible to obtain a high-quality solution, but the longer the computation time. In this work, the initial temperature is set to 200,000.
Inner-Loop Criterion: The inner loop simulates the process of the system reaching equilibrium at a given temperature. Theoretically, there should be a sufficient number of iterations. The more decision variables there are, the more iterations the inner loop should have. Therefore, in this paper, the inner loop stops when either the number of accepted solutions reaches $\lambda_{1}$ times of the number of variables, or the number of solutions generated exceeds $\lambda_{2}$ times of the number of variables. The value of $\lambda_{1},\lambda_{2}$ are tuned alongside other hyperparameters.
Outer-Loop Criterion and Cooling Rate: Generally, the termination temperature should be set to a sufficiently small positive number. The temperature update function follows a geometric rate: $T_{\text{curr}+1}=\alpha\cdot T_{\text{curr}}$, where $\alpha$ is set as 0.95 in this paper.
4 Experiments
We assess the capabilities of DCS in reducing class accuracy differences and enhancing overall accuracy with three ICL settings: 1-shot, 5-shot, and $N$-shot. Demonstrations in the prompt are randomly selected from the optimization set for 1-shot and 5-shot cases, while the $N$-shot demonstration cascades examples from each class, to minimize under-representations of any class. The proposed method is compared against SoTA methods, DNIP and FuRud. We prompt LLMs with an A100 GPU to obtain initial ICL outputs, following DNIP’s hyperparameters, then we run simulated annealing on CPUs. Evaluations are performed on seven multi-class text classification tasks across general and biomedical domains, including AGNews (AGN, 4-class;[38]), DBpedia (DBP, 14-class;[1]), SST-5 (SST, 5-class;[39]), TREC (6-class;[40, 41]), RTE (binary;[42]), DDI (5-class;[43]), and PubMedQA (PMQA, 3-class;[44]). Appendix C presents more details.
4.1 Main Results
Setup We perform quantitative evaluations on Llama-2-13B, for its representativeness as a widely applied open-source LLM. Indeed, Llama-2-13B is not particularly big in parameter size, but its architecture can be seen as basic building blocks for many advanced LLMs [45, 46, 47], and we expect similar findings in larger model variants or models similar in architecture, which will be discussed in Section 4.3. Table 1 reports our main findings.
Acc COBias
General Domain Biomedical Domain General Domain Biomedical Domain
Method All Avg AGN DBP SST TREC RTE Avg DDI PMQA All Avg AGN DBP SST TREC RTE Avg DDI PMQA
1-shot ICL + post-hoc debiasing
ICL 59.4 70.7 79.97.0 88.21.0 44.94.3 68.510.8 71.52.2 31.2 7.20.9 55.12.9 40.5 35.4 28.316.1 16.23.7 53.15.0 35.96.5 43.47.0 53.4 45.65.9 61.21.9
+DNIP 69.2 76.2 87.90.7 93.40.6 48.31.9 77.12.0 74.30.8 51.8 40.46.0 63.114.0 14.3 10.2 6.30.6 7.70.6 18.710.1 14.21.3 4.33.3 24.3 7.53.2 41.129.6
+FuRud 67.2 76.5 86.82.0 92.70.2 50.03.0 78.42.4 74.51.8 44.0 29.417.0 58.64.1 15.1 13.3 10.94.7 8.80.3 26.211.6 13.51.4 7.11.7 19.5 11.86.7 27.28.1
+DCS (Ours) 72.2 77.8 88.21.0 94.70.7 50.21.4 81.52.5 74.40.4 58.1 52.918.3 63.312.2 15.0 10.7 6.21.5 6.00.8 23.711.6 12.83.3 5.03.6 25.8 16.413.5 35.222.3
5-shot ICL + post-hoc debiasing
ICL 63.9 68.5 82.52.0 93.61.3 45.86.0 58.723.3 61.916.9 52.3 34.142.1 70.47.1 39.2 35.6 24.44.2 9.02.0 48.015.4 35.413.7 61.340.4 48.3 44.45.4 52.219.3
+DNIP 71.9 77.9 88.50.6 95.80.7 52.92.5 76.68.7 75.84.5 57.0 54.15.1 59.91.4 10.3 9.2 7.00.9 5.71.1 15.812.6 14.47.5 3.11.9 13.1 16.57.6 9.64.6
+FuRud 70.6 77.9 87.80.9 95.50.7 53.34.9 76.93.9 75.84.2 52.3 40.37.3 64.35.6 14.7 11.8 10.60.7 5.70.9 22.16.0 19.610.7 1.00.3 22.1 19.13.9 25.119.5
+DCS (Ours) 74.8 79.0 88.80.8 96.30.6 54.22.9 79.74.9 76.14.3 64.3 57.14.2 71.48.5 14.7 7.7 6.70.7 4.71.1 11.78.3 14.04.1 1.60.6 32.2 21.04.4 43.423.4
N-shot ICL + post-hoc debiasing
ICL 61.9 74.2 83.51.5 95.21.2 50.32.3 67.012.7 75.00.8 31.0 9.71.0 52.35.3 25.6 23.8 14.95.1 7.02.2 36.37.2 38.25.1 22.513.2 30.3 39.73.5 20.94.2
+DNIP 70.7 79.2 88.70.5 96.60.5 51.31.0 82.71.4 76.73.7 49.5 43.66.1 55.32.6 7.5 6.3 7.30.5 4.30.7 2.81.5 12.15.5 5.03.3 10.6 12.54.3 8.71.2
+FuRud 73.6 78.7 87.90.8 96.50.5 53.72.5 78.35.0 77.13.2 60.7 61.76.8 59.77.2 16.2 13.1 8.53.2 4.90.7 29.011.6 15.83.1 7.34.9 23.9 28.93.8 18.814.9
+DCS (Ours) 74.7 80.0 88.90.5 96.80.4 54.20.6 83.43.1 76.73.7 61.4 62.416.5 60.35.3 13.0 8.8 7.11.1 4.10.6 16.95.3 11.15.3 5.03.3 23.3 28.69.9 17.96.6
Table 1: Main results
The proposed method achieves SoTA overall accuracy improvements on all tasks except for RTE across all ICL settings. As for RTE, due to its simplicity as a binary classification task, all three methods obtain similar accuracy and COBias improvements. Meanwhile, it achieves SoTA COBias performance on general domain tasks. In general, its COBias is on par with the current best methods, with average COBias over all seven tasks falling between that of DNIP and FuRud. On AGNews, DBpeida, and TREC, the proposed method achieves both more overall accuracy improvements and more COBias reduction than DNIP and FuRud across all ICL settings, demonstrating the superior performance gains achieved by ensembling class-level and sample-level debiasing. Weak classes are elevated. We show in Section 4.2 that DCS powers up the low-accuracy classes. We observe that weaker classes are often corrected by rules, while stronger classes are most likely corrected by weights, further highlighting the need of an ensemble method.
Annealing Time
Figure 3: Task annealing time
Figure 3 plots the time used to optimize correction functions (left axis) and the search space size (right axis) for each task. For fair comparisons, we fix the weight scale for weight correction to 30 alongside 19 triangular membership functions, making a total of 49 possible correction functions and $49\cdot N$ possible assignments across all variables. DCS is computationally efficient where all tasks finish optimization within 150 outer loops, ranging from a maximum of 96.6 seconds to a minimum of 5.4 seconds. Notably, SST-5, DDI, and AGNews have a relatively longer time due to having a larger optimization set (Appendix C). In general, although larger optimization sets and search spaces need more time to converge, the annealing steps needed for a task can be further tuned to achieve cost-effective gains on downstream applications.
4.2 Overall Probability Correction Scheme
The overall probability correction scheme demonstrates the ratio of class-level corrections to sample-level corrections among all classes. Which type of correction should be used for a class is automatically determined by our method. Recall that: Class-level correction: Probabilities of the same output class are corrected uniformly using a same correction weight across all samples. It operates at the class level, not accounting for sample-based differences. Sample-level correction: This is a more refined correction using membership correction functions. Instead of a single weight, a unique mapping function is applied for each class. This allows varying degrees of correction across different samples - some probabilities may be adjusted significantly, others minimally, and some may remain unchanged - capturing the nuances of individual samples, leading to more precise adjustments. By combining them, our method ensures a balanced and adaptive correction, enhancing performances.
Dataset # Classes
# Classes that
use weight
corrections
# Classes that
use membership
corrections
Ratio of memberhsip
corrections to weight
corrections
#Classes that
use membership
corrections
in seed0/1/2 run
Remarks on any weakest class
that use membership
corrections
AGN 4 3.7 0.3 0.08 0/1/0
The weakest class Sports from seed1 run,
Acc: 0.75 (ICL) $\rightarrow$ 0.98 (corrected)
DBP 14 12.3 1.7 0.14 2/2/1 -
SST 5 4.3 0.7 0.16 1/0/1
The weakest class Terrible from seed0 run,
Acc: 0.21 (ICL) $\rightarrow$ 0.54 (corrected)
The weakest class Terrible from seed2 run,
Acc: 0.09 (ICL) $\rightarrow$ 0.64 (corrected)
TREC 6 5 1 0.20 0/2/1
The weakest class Entity from seed2 run,
Acc: 0.04 (ICL) $\rightarrow$ 0.76 (corrected)
RTE 2 0.7 1.3 1.86 1/2/1
The weakest class False from seed1 run,
Acc: 0.54 (ICL) $\rightarrow$ 0.78 (corrected)
DDI 5 2.7 2.3 0.85 3/2/2
The weakest class Negative from seed0 run,
Acc: 0 (ICL) $\rightarrow$ 0.73 (corrected)
The weakest class Negative from seed1 run,
Acc: 0 (ICL) $\rightarrow$ 0.46 (corrected)
The weakest class Negative from seed2 run,
Acc: 0.02 (ICL) $\rightarrow$ 0.93 (corrected)
PMQA 3 2.7 0.3 0.11 0/1/0
The weakest class Yes from seed1 run,
Acc: 0.38 (ICL) $\rightarrow$ 0.62 (corrected)
Table 2: Overall probability correction scheme ($N$-shot)
We exemplify the overall probability correction scheme with the $N$-shot case in Table 2. Numbers of classes are averaged over three runs of different demonstrations. As expected, most classes select class-level weight corrections. Membership corrections are less selected, but weakest classes benefit greatly from them, where 6 out of 7 tasks have selected membership corrections to correct their weakest classes (at least once over three runs). This demonstrates the necessity of enabling sample-level corrections to boost the performance of weak classes, and letting the optimization decide if sample-level corrections are needed.
4.3 Larger Models on DDI (Biomedical-domain)
Class Acc.
Model Acc. CB Prompt Demon. Neg. Eff. Mech. Adv. Int.
Llama-2-13B
seed0 0.0 0.0 100.0 0.0 2.3
seed1 0.0 91.4 0.3 17.7 39.6
ICL 7.2 45.6 seed2 0.0 0.8 84.1 54.8 88.5
seed0 32.5 38.1 32.1 29.0 32.3
seed1 62.1 30.6 32.8 42.1 31.3
+DCS 52.9 (+45.7) 16.4 (-29.2) seed2 80.7 12.8 8.6 6.3 11.5
Llama-2-70B
seed0 0.1 86.9 3.0 4.1 19.8
seed1 8.7 64.2 3.0 18.1 56.3
ICL 8.8 36.0 seed2 0.5 64.4 6.6 37.6 58.3
seed0 69.9 34.4 19.9 18.6 47.9
seed1 63.6 23.6 39.4 29.0 35.4
+DCS 63.7 (+54.9) 24.3 (-11.7) seed2 80.5 16.1 26.2 18.6 33.3
Table 3: Significant DDI Performance gains on both Llama-2-13B and Llama-2-70B; overall acc. and COBias use average scores over three runs.
Similar to its smaller variant Llama-2-13B, 1-shot ICL on Llama-2-70B also manifests class accuracy imbalances. Especially on the biomedical domain DDI classification, using 1-shot prompting, Llama-2-13B’s average ICL accuracy on DDI is 7.2% with COBias 45.6% as shown in Table 1, while Llama-2-70B obtains 8.8% accuracy with COBias 36.0% as shown in Table 3, showing that the larger variant is still highly prone to biased predictions on biomedical tasks. This suggests that the class accuracy imbalance may not be effectively alleviated by increasing model sizes, and the proposed post-hoc corrections can exactly help with the situation.
Due to effectively correcting probabilities for low-accuracy classes as shown by Table 3, our method demonstrates significant accuracy enhancements on DDI, across both small and large model variants. For more Llama-2-70B performance gains, detailed scores can be found in Appendix D.
4.4 Ablations
[…]
AGN DBP SST TREC RTE DDI PMQA
Prompt Obj. Acc. CB. Acc. CB. Acc. CB. Acc. CB. Acc. CB. Acc. CB. Acc. CB.
Full 88.2 6.2 94.7($\uparrow$6.1) 6.0($\downarrow$10.2) 50.2 23.7 81.5($\uparrow$12.9) 12.8($\downarrow$23.1) 74.4 5.0 52.9 16.4 63.3 35.2
ZAcc 88.4 7.8 92.8($\uparrow$4.2) 8.6($\downarrow$7.6) 51.4 50.9 80.6($\uparrow$12.0) 16.7($\downarrow$19.2) 75.7 12.8 82.2 39.8 71.3 58.8
1-shot ZAcc+ZPMI 88.5 7.5 94.1($\uparrow$5.5) 7.5($\downarrow$8.7) 51.7 53.0 81.1($\uparrow$12.5) 15.1($\downarrow$20.8) 75.8 13.0 82.9 40.0 71.1 59.1
Full 88.8 6.7 96.3($\uparrow$2.7) 4.7($\downarrow$4.3) 54.2 11.7 79.7($\uparrow$21.0) 14.0($\downarrow$21.4) 76.1 1.6 57.1 21.0 71.4 43.4
ZAcc 88.9 7.8 95.0($\uparrow$1.4) 7.1($\downarrow$1.9) 54.2 41.6 79.5($\uparrow$20.8) 14.6($\downarrow$20.8) 76.2 13.7 82.8 40.0 76.7 58.8
5-shot ZAcc+ZPMI 88.9 7.9 95.9($\uparrow$2.3) 5.5($\downarrow$3.5) 54.9 41.4 78.9($\uparrow$20.2) 21.0($\downarrow$14.4) 76.9 10.2 82.9 40.2 76.8 58.8
Full 88.9 7.1 96.8($\uparrow$1.7) 4.1($\downarrow$2.8) 54.2 16.9 83.4($\uparrow$16.4) 11.1($\downarrow$27.1) 76.7 5.0 62.4 28.6 60.3 17.9
ZAcc 88.8 7.5 96.3($\uparrow$1.2) 5.1($\downarrow$1.8) 54.4 37.4 81.9($\uparrow$14.9) 18.4($\downarrow$19.8) 76.9 17.0 82.8 40.1 76.0 60.5
N-shot ZAcc+ZPMI 89.0 7.5 96.5($\uparrow$1.4) 4.8($\downarrow$2.1) 54.8 40.8 83.3($\uparrow$16.3) 15.3($\downarrow$22.9) 77.5 15.7 82.9 40.1 76.1 60.7
Table 4: ZAcc objective ablations; average scores over three random seeds are reported. Improvements over ICL scores are shown by ($\uparrow$) and ($\downarrow$).
[…]
5 Conclusion and Limitations
This paper proposes DCS, a post-hoc debiasing framework that unifies sample and class level corrections to enrich low-accuracy classes in ICL outputs. It achieves SoTA accuracy on seven benchmark text classification tasks, while effectively mitigating class performance imbalances. The resulting probability correction scheme clearly shows that combining membership correction and weight correction is essential to elevate weak classes. For limitations and future work, the post-hoc correction framework can be extended to other useful scenarios, such as text generation. Combining DCS into the decoding process can help rectify the probabilities of top candidates for the first token (or first several tokens) to lead a better, less biased response. More future directions are discussed in Appendix E.
Appendix A Discussion on Solution Methods for NIP
Typical solutions for combinatorial optimization problems include the branch-and-bound (B&B) algorithm, which is useful for solving mixed-integer linear programs [48]. Solvers are made into both open-source software such as CVXPY [49, 50] or SCIP [51], and commercial solvers like Gurobi [52] or LINGO [53].
We need to point out that B&B and other related algorithms are not very suitable to search solutions for the proposed nonlinear integer programming based optimization model, which contain discontinuous functions and is inherently non-differentiable. Relaxation to subproblems is needed if B&B has to be used. Relaxation can be done by creating Lagrange multipliers and Lagrangian functions for each constraint, which is often easier to solve than the original nonlinear integer programming problem, because only continuous variables are involved [54, 55]. The relaxed problem finds the optimal value of the Lagrangian function, which is a lower bound for the original problem, and this optimal value of the relaxed problem is used as a bound in the B&B search. However, solving the subproblems efficiently becomes a challenge, as B&B is generally not a polynomial-time algorithm.
Instead, we can opt for metaheuristics [56], such as simulated annealing [35, 36], to find the optimal solution. Although heuristic search can get slower with increasing search space size [57], its computational complexity is within polynomial time [5]. To enable application of our proposed method to tasks with even more classes, we adopt simulated annealing (SA) to solve our model.
In SA, we define an initial temperature $T$ and an annealing schedule including the cooling rate and the stopping criteria. In each iteration, a neighborhood solution is generated by perturbing the current solution. The algorithm calculates the objective function value for the new solution (Step Evaluate Z in Fig. 2). If the new solution is better, accept it; if the new solution is worse, accept it with a probability determined by the Metropolis criterion. Then the temperature is reduced according to the cooling rate, until convergence. In addition, the initial solution can also be randomly generated, where $N$ numbers are randomly selected from $\mathbb{D}=\{1,\dots,D_{F}+D_{W}\}$ as the initial values for the selection variables $\xi_{j}$’s.
Appendix B List of notations
The parameters, sets and decision variables used throughout this paper are listed in Table 5.
Symbol Description
Parameters
$p_{m}$ The ICL output token’s class probabilities of the $m$-th instance in a dataset
$\omega_{k}$ The $k$-th weight coefficient where $k\in\{1,...,D_{F}\}$
$\mu_{k}$ The $k$-th membership function where $k\in\{D_{F}+1,...,D_{F}+D_{W}\}$
Sets
$\mathbb{R}^{\text{Weight}}$ The set of $D_{W}$ weight coefficients, equally spaced discrete values from 0 to 1
$\mathbb{R}^{\text{Membership}}$ The set of $D_{F}$ triangular membership functions
$\mathbb{D}$ The set of selection indices
Variables
$y_{m},\hat{y}_{m}$ The label and prediction of the $m$-th instance
$\xi_{j}$ The integer selection variable for class $j$
Table 5: List of notations
Appendix C Experimental Setups
For each evaluation task, we tune the weight scale for weight correction, $\beta$ and $\tau$ used in the objective function, and the simulated annealing algorithm’s number of outer loops, inner loops, the initial temperature, and stopping criteria (stopping temperature, the maximal number of solutions generated, and the maximal number of accepted solutions) on a development set. The optimization/development set split is 0.95/0.05. Dataset preprocessing follows DNIP [5], and their statistics are shown in Table 6.
Dataset
Optimization
set size
(incl. dev)
Evaluation
set size
Prompt
AGN 10,000 5,000
Please classify the news articles into the categories of
[…]
Article: [Input]
Answer:
DBP 10,000 5,000
Please classify the documents based on whether they are
about a Company, School, Artist, Athlete, Politician,
Transportation, Building, Nature, Village, Animal,
[…]
Article: [Input]
Answer:
SST 8,544 2,210
Review: [Input]
Sentiment:
TREC 5,452 500
Please classify the questions based on whether their answer
type is a Number, Location, Person, Description, Entity,
[…]
Question: [Input]
Answer Type:
RTE 2,490 277
[Input premise]
question:
[Input hypothesis]
True or False?
answer:
DDI 10,000 5,716
Please choose a most suitable answer from Negative, Effect,
Mechanism, Advice, or Interaction, for the drug-drug
interaction relation between the @drug\$ pair in the
[…]
Description: [Input]
Answer:
PMQA 1,000 500
Please choose a most suitable answer from yes, no, or maybe,
[…]
Context: [Input context]
Question: [Input question] yes, no, or maybe?
Answer:
Table 6: Dataset statistics
Appendix D Full Performance Comparisons on Llama-2-70B
Our method effectively applies to larger LLM variants, showcased by Llama-2-70B. Debiasing results using DCS on 1-shot ICL outputs in Table 7 show consistent improvements in average overall accuracy across three seeds of demonstrations, outperforming DNIP and FuRud. Our method achieves better COBias reduction than the other two methods on general domain tasks, and
Acc COBias
General Domain Biomedical Domain General Domain Biomedical Domain
Method All Avg AGN DBP SST TREC RTE Avg DDI PMQA All Avg AGN DBP SST TREC RTE Avg DDI PMQA
1-shot ICL + post-hoc debiasing
ICL 66.4 76.2 87.42.2 94.11.0 41.88.6 79.32.5 78.62.6 41.7 8.83.7 74.52.3 31.8 25.1 14.55.6 8.31.2 54.71.4 22.42.0 25.57.6 48.8 36.02.1 61.53.3
+DNIP 73.6 79.0 89.80.2 96.50.5 49.42.0 79.34.0 80.01.1 60.0 52.514.4 67.52.4 17.3 15.9 6.50.4 3.80.3 47.114.0 17.34.8 4.61.8 21.1 16.49.0 25.72.0
+FuRud 75.9 79.0 89.40.6 96.20.8 50.02.2 79.53.8 79.99.1 68.1 60.011.8 76.12.2 21.4 15.0 7.11.0 3.80.6 40.18.3 19.05.2 4.81.5 37.5 25.27.5 49.79.4
+DCS 76.8 79.6 89.90.2 96.70.5 50.11.7 81.03.6 80.11.1 70.0 63.76.2 76.22.4 20.7 14.3 6.60.1 3.60.2 40.38.2 15.65.6 5.31.6 36.8 24.35.5 49.29.1
Table 7: Quantitative evaluations on Llama-2-70B
The takeaway is, increasing LLM sizes without post-hoc corrections may not always be an effective way to resolve output class accuracy imbalances, which further validates the necessity of the proposed post-hoc correction method.
Appendix E More Future Directions
This work deals with prediction biases when directly prompting LLMs. When LLMs or smaller LMs like BERT [58] are finetuned on tasks, which may result in further imbalanced class accuracies, post-hoc corrections can make class accuracies fairer. It would be valuable to understand how these output class probabilities differ from those by direct prompting, and enable custom extensions to the current framework to provide more targeted corrections. It would also be interesting to see how it may help with data-filtered or data-augmented finetuning. In terms of interpretability, those classes that apply class-level correction to any instance is interpretable only at the broader level, so further investigations on how to enforce more interpretability to weight corrections could be conducted.